OpenStudy (calculusxy):

the sum of five consecutive multiples of 7 is 350. the smallest of these is a. 28 b. 35 c. 42 d. 56 e. 70

OpenStudy (texaschic101):

x + (x + 7) + (x + 14) + (x + 21) + (x + 28) = 350 -- combine like terms 5x + 70 = 350 5x = 350 - 70 5x = 280 x = 56 x + 7 = 56 + 7 = 63 x + 14 = 56 + 14 = 70 x + 21 = 56 + 21 = 77 x + 28 = 56 + 28 = 84 check.. 56 + 63 + 70 + 77 + 84 = 350 350 = 350 (correct) your lowest number (x) = 56 any questions ?

OpenStudy (calculusxy):

i don't have any questions. thanks a lot @texaschic101

OpenStudy (texaschic101):

no problem :)